OpenStudy (tylerd):

he function f(x)=ln(9−x) is represented as a power series

OpenStudy (tylerd):

\[f(x)=\sum_{n=0}^{oo}c_{n}x^n\] find c0, c1,c2 etc.

OpenStudy (tylerd):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

check out this page https://www.math.ucdavis.edu/~thomases/W11_16C1_lec_3_4&3_7_11.pdf specifically what I'm attaching as an image

OpenStudy (tylerd):

ok thanks, the book i have is more confusing then this

jimthompson5910 (jim_thompson5910):

np

OpenStudy (tylerd):

how do i solve for c?

OpenStudy (tylerd):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

c is where you center the power series at

jimthompson5910 (jim_thompson5910):

so if you center it at say x = 0, then c = 0 or if you center it at x = 1, then c = 1

OpenStudy (tylerd):

im trying to find f'(c)(x-c) so i get 1/9(x-0) = x/9 but what is x?

jimthompson5910 (jim_thompson5910):

can I see the full problem as a screenshot?

jimthompson5910 (jim_thompson5910):

on that pdf, the first term is f(c) since the first term you have is ln(9) and f(c) = ln(9), this means f(x) = ln(9-x) f(c) = ln(9-c) ln(9) = ln(9-c) 9 = 9-c c = 0

jimthompson5910 (jim_thompson5910):

so this power series is centered at x = 0

OpenStudy (tylerd):

yes i got that far

OpenStudy (tylerd):

but the next one c1

OpenStudy (tylerd):

c1 is not = 0

jimthompson5910 (jim_thompson5910):

c1 = f ' (c) c1 = -(9-c)^(-1) c1 = -(9-0)^(-1) c1 = -1/9

jimthompson5910 (jim_thompson5910):

f ' (x) = -(9-x)^(-1)

jimthompson5910 (jim_thompson5910):

c2 = f '' (c)/2 c3 = f ''' (c)/6 etc etc look at page 5 of that pdf

OpenStudy (tylerd):

how did you get negative?

jimthompson5910 (jim_thompson5910):

the derivative of ln(x) is 1/x so the derivative of ln(9-x) is 1/(9-x) times the derivative of the inside terms. Don't forget the chain rule

jimthompson5910 (jim_thompson5910):

1/(9-x) * derivative of (9-x) = (9-x)^(-1) * (-1) = -(9-x)^(-1)

OpenStudy (tylerd):

ah ok

OpenStudy (tylerd):

so x and c are the same

OpenStudy (tylerd):

in most of these problems or?

jimthompson5910 (jim_thompson5910):

the same? what do you mean

jimthompson5910 (jim_thompson5910):

x is a variable, c is a constant

OpenStudy (tylerd):

c=0 f(0)=f'(0)(x-0)?

jimthompson5910 (jim_thompson5910):

so it looks like you're computing the first term

OpenStudy (tylerd):

yes

OpenStudy (tylerd):

-1/9(x)

OpenStudy (tylerd):

ah ok so -1/9 is just the c part

jimthompson5910 (jim_thompson5910):

yes the coefficient to the (x-c)^n term

OpenStudy (tylerd):

alright i should be able to get the rest thanks alot man.

OpenStudy (tylerd):

life saver

jimthompson5910 (jim_thompson5910):

glad to be of help